### v2

# Multi-Modal and Multi-Agent Systems Meet Rationality: A Survey

Abstract
Rationality is the quality of being guided by reason, characterized by logical thinking and decision-making that align with evidence and logical rules. This quality is essential for effective problem-solving, as it ensures that solutions are well-founded and systematically derived. Despite the advancements of large language models (LLMs) in generating human-like text with remarkable accuracy, they present biases inherited from the training data, inconsistency across different contexts, and difficulty understanding complex scenarios involving multiple layers of context. Therefore, recent research attempts to leverage the strength of multiple agents working collaboratively with various types of data and tools for enhanced consistency and reliability. To that end, this paper aims to understand whether multi-modal and multi-agent systems are advancing toward rationality by surveying the state-of-the-art works, identifying advancements over single-agent and single-modal systems in terms of rationality, and discussing open problems and future directions. We maintain an open repository at https://github.com/bowen-upenn/MMMA_Rationality.
Multi-Modal and Multi-Agent Systems Meet Rationality: A Survey
Bowen Jiang1, Yangxinyu Xie1, 2, Xiaomeng Wang1, Weijie J. Su1 Camillo J. Taylor1, Tanwi Mallick2    University of Pennsylvania1    Argonne National Laboratory2 Philadelphia, PA, 19104, USA   Lemont, IL, 60439, USA {bwjiang@seas, xinyux@wharton, cjtaylor@seas}.upenn.edu, tmallick@anl.gov
1 Introduction
Large language models (LLMs) have demonstrated promising results across a broad spectrum of tasks, particularly in exhibiting capabilities that plausibly mimic human-like reasoning Wei et al. (2022); Yao et al. (2024); Besta et al. (2024); Shinn et al. (2024); Bubeck et al. (2023); Valmeekam et al. (2023); Prasad et al. (2023). These models leverage the richness of human language to abstract concepts, elaborate thinking process, comprehend complex user queries, and develop plans and solutions in decision-making scenarios. Despite these advances, recent research has revealed that even state-of-the-art LLMs exhibit various forms of irrational behaviors, such as the framing effect, certainty effect, overweighting bias, and conjunction fallacy Binz and Schulz (2023); Echterhoff et al. (2024); Mukherjee and Chang (2024); Macmillan-Scott and Musolesi (2024); Wang et al. (2024a); Suri et al. (2024). These biases significantly challenge the utility of LLMs in natural language processing research. For example, LLM-based evaluators, a popular choice for automated assessments for text generation, display cognitive biases against certain responses irrespective of their actual quality or relevance Stureborg et al. (2024); Koo et al. (2023). Irrationality and hallucinations  Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023) also undermine the practical deployment of LLMs in critical sectors like healthcare, finance, and legal services He et al. (2023); Li et al. (2023i); Kang and Liu (2023); Cheong et al. (2024), where reliability and consistency are paramount. The emerging concern about the factual accuracy and trustworthiness of LLMs highlighting an urgent need to develop better agents or agent systems Nakajima (2023); Gravitas (2023) with rational reasoning processes.
One possible reason for the LLMs’ irrational behaviors, as suggested by Bubeck et al. (2023) and Sun (2024), is the autoregressive nature of existing language models. This architecture doesn’t allow for an “internal scratchpad" beyond these models’ inner parametric representations of knowledge, causing them to fail to reason rationally when faced with problems that require more complex and iterative procedures. Thus, an important question emerges: How can we design an LLM-based agent capable of rational decision-making that can overcome these biases and inconsistencies?
Recent advancements in multi-modal and multi-agent frameworks offer a promising direction to address this challenge, which leverage the expertise of different agents acting together towards a collective goal. Multi-modal foundation models Awadalla et al. (2023); Liu et al. (2023b); Wang et al. (2023d); OpenAI (2023); Reid et al. (2024) enhance reasoning by grounding decisions in a broader sensory context, akin to how human brains integrate rich sensory inputs to form a more holistic base of knowledge. Meanwhile, multi-agent systems introduce mechanisms such as consensus, debate, and self-consistency Du et al. (2023); Liang et al. (2023); Talebirad and Nadiri (2023); Madaan et al. (2024); Cohen et al. (2023); Shinn et al. (2024); Mohtashami et al. (2023), which allow for more refined and reliable output through collaborative interaction among multiple instances. Each agent is specialized in different domains and offers its unique perspectives, simulating the dynamics of discussion in human societies. Multi-agent systems can also incorporate multi-modal agents and agents specialized in querying external knowledge sources or tools Lewis et al. (2020); Schick et al. (2024); Tang et al. (2023); Pan et al. (2024) to overcome hallucinations, ensuring that their results are more robust, deterministic, and trustworthy, thus significantly improving the quality of the generated responses towards rationality.
This survey provides a unique lens to interpret the underlying motivations behind current multi-modal and/or multi-agent systems. Drawing from cognitive science, we first delineate four fundamental requirements for rational thinking in Section 2, and then Section 4 discusses how research fields within the multi-modality and multi-agents literature are progressing towards rationality by inherently improving these criteria. We posit that such advancements are bridging the gap between the performance of these systems and the expectations for a rational thinker, in contrast to traditional single-agent language-only models. Furthermore, Section 5 highlights the lack of sufficient evaluation metrics and benchmarks in the existing literature to adequately measure the rationality of LLMs or agent systems. We hope this survey can inspire further research at the intersection between agent systems and cognitive science.
2 Defining Rationality
Figure 1: The evolutionary tree of multi-agent and/or multi-modal systems related to the four axioms of rationality. Many proposed approaches strive to address multiple axioms simultaneously. Bold fonts are used to mark works that involve multi-modalities. This tree also includes foundational works to provide a clearer reference of time.
A rational agent should avoid reaching contradictory conclusions in decision making processes, respecting the physical and factual reality of the world in which it operates. Thus, drawing on foundational works in rational decision-making (Tversky and Kahneman, 1988; Hastie and Dawes, 2009; Eisenführ et al., 2010), this section adopts an axiomatic approach to define rationality, presenting four substantive axioms that we expect a rational agent or agent systems to fulfill:
Grounding
The decision of a rational agent is grounded on the physical and factual reality. For example, a flight booking agent must accurately retrieve available airports without fabricating nonexistent ones, and a video generation agent should adhere to the laws of physics in a world model.
Naturally, in order to make a sound decision, the agent must be able to integrate sufficient and accurate information from different sources and modalities grounded in reality without hallucination. While this requirement is generally not explicitly stated in the cognitive science literature when defining rationality, it is implicitly implied, as most humans have access to physical reality through multiple sensory signals.
Orderability of Preferences
When comparing alternatives in a decision scenario, a rational agent can rank the options based on the current state and ultimately select the most preferred one based on the expected outcomes. This orderability consists of several key principles, including comparability, transitivity closure, solvability, etc. with detailed defined in Appexdix A. The orderability of preferences ensures the agent can make consistent and logical choices when faced with multiple alternatives. LLM-based evaluations heavily rely on this property, as discussed in Appendix B.
Independence from irrelevant context
The agent’s preference should not be influenced by information irrelevant to the decision problem at hand. LLMs have been shown to exhibit irrational behavior when presented with irrelevant context Shi et al. (2023); Wu et al. (2024); Liu et al. (2024d); Yoran et al. (2023), leading to confusion and suboptimal decisions. To ensure rationality, an agent must be able to identify and disregard irrelevant information, focusing solely on the factors that directly impact the decision-making processes.
Invariance
The preference of a rational agent remains invariant across equivalent representations of the decision problem, regardless of specific wordings or modalities.
3 Scope
Unlike existing surveys Han et al. (2024); Guo et al. (2024); Xie et al. (2024a); Durante et al. (2024); Cui et al. (2024); Xu et al. (2024b); Zhang et al. (2024a); Cheng et al. (2024); Li et al. (2024a) that focus on the components, structures, agent profiling, planning, communications, memories, and applications of multi-modal and/or multi-agent systems, this survey is the first to specifically examine the increasingly important relations between rationality and these multi-modal and multi-agent systems, exploring how they contribute to enhancing the rationality in decision-making processes.
We emphasize that rationality, by definition, is not equivalent to reasoning Khardon and Roth (1997); Huang and Chang (2022); Zhang et al. (2024a); Qiao et al. (2022), although deeply intertwined. Rationality involves making logically consistent decisions grounded with reality, while reasoning refers to the cognitive process of drawing logical inferences and conclusions from available information, as illustrated in the following thought experiment:
Consider an environment where the input space and the output decision space are finite. A lookup table with consistent mapping from input to output is inherently rational, while no reasoning is necessarily present in the mapping.
Despite this example, it is still crucial to acknowledge that reasoning typically plays a vital role in ensuring rationality, especially in complex and dynamic real-world scenarios where a simple lookup table is insufficient. Agents must possess the ability to reason through novel situations, adapt to changing circumstances, make plans, and achieve rational decisions based on incomplete or uncertain information.
Rationality is also different from Theory of Mind (ToM) Apperly and Butterfill (2009); Nye et al. (2021); Oguntola et al. (2021); Hagendorff (2023); Li et al. (2023c); Sclar et al. (2023b); Kosinski (2023) in machine psychology. ToM refers to the model’s ability to understand that others’ mental states, beliefs, desires, emotions, and intentions may be different from its own.
4 Towards Rationality through Multi-Modal and Multi-Agent Systems
This section surveys recent advancements in multi-modal and multi-agent systems, categorized by their fields as depicted in Figure 1.
Each category of research in Figure 1, such as knowledge retrieval or neuro-symbolic reasoning, addresses one or more fundamental requirements for rational thinking. These rationality requirements are typically intertwined; therefore, an approach that enhances one aspect of rationality often inherently improves others simultaneously. Meanwhile, the overall goal of current multi-agent system in achieving rationality can usually be distilled into two key concepts: deliberation and abstraction. Deliberation encourages slower reasoning process such as brainstorming and reflection, while abstraction refers to boiling down the problem its into logical essence like calling APIs of tools or incorporating neuro-symbolic reasoning agents.
Most existing studies do not explicitly base their frameworks on rationality in their original writings. Our analysis aims to reinterpret these works through the lens of our four axioms of rationality, offering a novel perspective that bridges existing methodologies with rational principles.
Besides, dual-process theories Sun (2001); Evans (2003); Kahneman (2011); Sun (2024); Nye et al. (2021) in human cognition suggest that a single LLM, in analogy, might engage primarily in “System 1" thinking that is fast and automatic, but prone to biases. In contrast, “System 2" thinking is slow and rule-based (Kahneman, 2011), which is more reliable but requires more cognitive effort. This section will also discuss how multi-modal and multi-agent systems align with these theories.
4.1 Towards Grounding & Invariance through Multi-Modal Models
Multi-modal approaches aim to improve the information grounding across various channels, such as language and vision. By incorporating multi-modal agents, multi-agent systems can greatly expand their capabilities, enabling a richer, more accurate, and contextually aware interpretation of environment.
Multi-Modal Foundation Models
Grounding an agent solely based on textual language can be challenging, as information can be represented much more efficiently through other sensory modes. For example, as a picture is worth a thousand words, recent advances in large vision-language pretraining have enabled LLMs with robust language comprehension capabilities to finally perceive the visual world. Multi-modal foundation models, including but not limited to CLIP Radford et al. (2021), VLBERT and ViLBERT Su et al. (2019); Lu et al. (2019), BLIP-2 Li et al. (2023e), (Open) Flamingo Alayrac et al. (2022); Awadalla et al. (2023), LLaVA Liu et al. (2024c, 2023b), CogVLM Wang et al. (2023d), MiniGPT-4 Zhu et al. (2023a), GPT-4 Vision OpenAI (2023) and GPT-4o OpenAI (2024), and Gemini 1.5 Pro Reid et al. (2024) serve as the cornerstones for multi-modal agent systems to ground knowledge in vision and beyond.
Alignment with Dual-Process Theories
Generating output texts from input images in these models requires only a single inference pass, which is quick and straightforward. This process bypasses the need for iterative reasoning or reflective steps, aligning single-model inference closely with the System 1 process of fast and automatic thinking.
The adaptation of Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022), a technique popularized in language-only models, in LLaVA-RLHF Sun et al. (2023b) demonstrates promising advancements in reducing hallucination from cross-modal misalignment. Furthermore, visual instruction-tuning Liu et al. (2024c); Dai et al. (2024); Bai et al. (2023); Wang et al. (2023c) enables advanced foundation models like LLaVA Liu et al. (2024c, 2023b), GPT-4 Vision OpenAI (2023), and Gemini 1.5 Pro Reid et al. (2024) to engage in more detailed instruction following, multi-round human-agent interactions, and collaborations with other agents, performing deliberate multimodal reasoning, tool using Schick et al. (2024), and strategic thinking to understand more engineered or context-dependent queries. This opens the possibility of subsequent research on the System 2 process in multi-modal foundation models.
Information Grounding
Multi-modalities help enhance the functionality of agent systems through more diverse information grounding. For example, Chain-of-Action Pan et al. (2024) advances the single-modal Search-in-the-Chain Xu et al. (2023) by supporting multi-modal data retrieval for faithful question answering. DoraemonGPT Yang et al. (2024) decomposes complex tasks into simpler ones toward understanding dynamic scenes, where multi-modal understanding is necessary for spatial-temporal videos analysis. RA-CM3 Yasunaga et al. (2022) augments baseline retrieval-augmented LLMs with raw multi-modal documents that include both images and texts, assuming that these two modalities can contextualize each other and make the documents more informative, leading to better generator performance. The multi-modal capabilities also allow HuggingGPT Shen et al. (2024b), Agent LUMOS Yin et al. (2023), ToolAlpaca Tang et al. (2023), and AssistGPT Gao et al. (2023b) to expand the scope of tasks they can address, including cooperation among specialized agents or tools capable of handling different information modalities.
Web agents are another example of how multi-modal agents surpass language-only ones. In agents like Pix2Act Shaw et al. (2024), WebGUM Furuta et al. (2023), CogAgentHong et al. (2023b), and SeeAct Zheng et al. (2024a), web navigation is grounded on graphical user interface (GUI) rather than solely on HTML texts Shen et al. (2024a); Yao et al. (2022a); Deng et al. (2024); Gur et al. (2023). This method of visual grounding offers higher information density compared to HTML codes that are usually lengthy, noisy, and sometimes even incomplete Zheng et al. (2024a). Supporting the importance of vision, ablation studies in WebGUM Furuta et al. (2023) also reports $5.5\%$ success rate improvement on the MiniWoB++ dataset Shi et al. (2017); Liu et al. (2018) by simply adding the image modality.
Large world models is an emerging and promising direction to reduce multimodal hallucinations. The notion is also mentioned in “Objective-driven AI" LeCun (2024), where agents have behavior driven by fulfilling objectives, i.e., drives, and they understand how the world works with common sense knowledge, beyond an auto-regressive generation. LeCun (2024) proposes the urgency for agents to learn to reason beyond feed-forward, i.e., the System 1 subconscious computation, and start making System 2 reasoning and planning on complicated actions to satisfy objectives with a grounding on world models. For example, Ghost-in-the-Minecraft Zhu et al. (2023b) and Voyager Wang et al. (2023a) have agents living in a well-defined game-world environment. JEPA LeCun (2022) creates a recurrent world model in an abstract representation space. Large World Model (LWM) Liu et al. (2024b) and Sora Brooks et al. (2024) develop insights from both textual knowledge and the world through video sequences. They both advance toward general-purpose simulators of the world, but still lack reliable physical engines for guaranteed grounding in real-world dynamics.
Invariance Across Modalities
Achieving representation invariance across modalities is another critical facet of rationality. Given adequate information grounding, agents should make consistent decisions across different modalities that share equivalent underlying logic.
Multi-modal foundation models are particularly adept at promoting invariance by processing multi-modal data in an unified representation. Specifically, their large-scale cross-modal pretraining stage seamlessly tokenizes both vision and language inputs into a joint hidden embedding space, learning cross-modal correlations through a data-driven approach. In other words, image tokens are simply regarded as a foreign language Wang et al. (2022a). Moreover, the cross-modal validation inherent in multi-modal foundation models allows for reconciliation of data from different modalities, closing their distance in the hidden embedding space Radford et al. (2021).
The concept of invariance is the cornerstone of Visual Question Answering (VQA) agents Chen et al. (2022); Jiang et al. (2024); Wang et al. (2023e); Yi et al. (2018); Wang et al. (2022a); Bao et al. (2022); Zhao and Xu (2023). On one hand, these agents must grasp the invariant semantics of any open-ended questions posed about images, maintaining consistency despite variations in wording, syntax, or language. On the other hand, within a multi-agent VQA system, visual agents can provide crucial verification and support for language-based reasoning Wang et al. (2023e); Jiang et al. (2024); Zhao and Xu (2023), while language queries can direct the attention of visual agents, based on a shared and invariant underlying knowledge across vision and language domains.
4.2 Towards Grounding through Knowledge Retrieval
Bounded Rationality March and Simon (1958); Selten (1990) is a concept tailored to cognitively limited agents, suggesting that decision-making is limited by the resources available at hand, and any deviations from the optimal are primarily due to insufficient computational capacity and bounded working memory. In terms of LLMs, the parametric nature of their existing architecture Vaswani et al. (2017) fundamentally limits how much information they can hold. As a result, in the face of uncertainty, LLMs often hallucinate  Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023), generating outputs that are not supported by the factual reality of the environment. Retrieval-Augmented Generation (RAG) Lewis et al. (2020) marks a significant milestone in addressing such an inherent limitation of LLMs. Broadly speaking, RAG refers to any mechanism that provides external knowledge to the input context of an LLM and helps it deliver responses with up-to-date, factual, and grounded information.
A multi-modal and/or multi-agent system can include planning agents in its framework, which is akin to the System 2 process that can determine how and where to retrieve external knowledge, and what specific information to acquire. Additionally, the system can have summarizing agents that utilize retrieved knowledge to enrich the system’s language outputs with better factuality.
There are multiple works that construct large-scale knowledge graphs (KGs) Hogan et al. (2021) from real-world sources to effectively expand their working memory. Specifically, compared to language-only models, MAVEx Wu et al. (2022) improves system’s scores by $9.5\%$ compared to an image-only baseline through the integration of knowledge from ConceptNet Speer et al. (2017) and Wikipedia Wikipedia contributors (2004). It also improves the scores by $8.3\%$ by using the image modality for cross-modal validations with an oracle. Thanks to the external knowledge base, ReAct Yao et al. (2022b) reduces false positive rates from hallucination by $8.0\%$ compared to CoT Wei et al. (2022). CuriousLLM Yang and Zhu (2024) presents ablation studies showing the effectiveness of KGs on improving reasoning within the search process. MineDojo Fan et al. (2022) observes that internet-scale multi-modal knowledge allows models to significantly outperform all creative task baselines. Equipped with world knowledge, RA-CM3 Yasunaga et al. (2022) can finally generate faithful images from captions compared to CM3 Aghajanyan et al. (2022) and Stable Diffusion Rombach et al. (2022). CooperKGC Ye et al. (2023) enables multi-agent collaborations, leveraging knowledge bases of different experts. It finds that the incorporation of KGs improves F1 scores by $10.0$-$33.6\%$ across different backgrounds, and adding more collaboration rounds also enhance performance by about $10.0$-$30.0\%$. DoraemonGPT Yang et al. (2024) supports knowledge tools to assist the understanding of specialized video contents. SIRI Wang et al. (2023e) builds a multi-view knowledge base to increase the explainability of visual question answering. Grounding agents in external knowledge base also promotes more factual rationales and fewer hallucinations, especially in scientific and medical domains, exemplified by Chameleon Lu et al. (2024), Chain-of-Knowledge Li et al. (2023g), WildfireGPT Xie et al. (2024b), and Agent Hospital Li et al. (2024b). Chain-of-Knowledge Li et al. (2023g) even discovers that integrating multiple knowledge sources enhances performance by $2.1\%$ compared to using a single source in its experiments.
4.3 Towards Grounding & Invariance & Independence from Irrelevant Contexts through Tool Utilization
Information Grounding
Enabling agents to use tools also expands their bounded working memories, akin to retrieving external knowledge. Toolformer Schick et al. (2024) opens a new era that allows LLMs to use external tools via API calls following predefined syntax, effectively extending their capabilities beyond their intrinsic limitations and enforcing consistent and predictable outputs.
A multi-agent system can coordinate agents understanding when and which tool to use, which modality of information the tool should expect, how to call the corresponding API, and how to incorporate outputs from the API calls, which anchors subsequent reasoning processes with more accurate information beyond their parametric memory. For example, VisProg Gupta and Kembhavi (2023), ViperGPT Surís et al. (2023), and Parsel Zelikman et al. (2023) generate Python programs to reliably execute subroutines. Gupta and Kembhavi (2023); Surís et al. (2023) also invoke off-the-shelf models for multimodal assistance. Foundation models are not specifically trained for object detection or segmentation, so BuboGPT Zhao et al. (2023) and Multi-Agent VQA Jiang et al. (2024) call SAM Kirillov et al. (2023); Ren et al. (2024) as the tool. Jiang et al. (2024) finds $8.8\%$ of accuracy improvements compared to a single agent. Besides, BabyAGI Nakajima (2023), Chamelon Lu et al. (2024), AssistGPT Gao et al. (2023b), Avis Hu et al. (2024), ToolAlpaca Tang et al. (2023), MetaGPT Hong et al. (2023a), Agent LUMOS Yin et al. (2023), AutoAct Qiao et al. (2024), $\alpha$-UMi Shen et al. (2024a), and ConAgents Shi et al. (2024) harness compositional reasoning to enable generalized multi-agent systems with planning and modular tool-using capabilities in real-world scenarios.
Abstraction that Boils Down to Logical Essence
In most cases, tools require translating natural language queries into API calls with predefined syntax. Once the planning agent has determined the APIs and their input arguments, the original queries that may contain irrelevant contexts become invisible to the tools, and the tools will ignore any variance in the original queries as long as they share the equivalent underlying logic. Take Multi-Agent VQA Jiang et al. (2024) as an example. It has an LLM which provides only relevant object names rather than the whole visual question to the Grounded SAM Ren et al. (2024) component of the system acting as an object-detector. Similarly, the image editing tools in VisProg Gupta and Kembhavi (2023) only receive a fixed set of arguments translated from user queries to perform deterministic code executions. SeeAct Zheng et al. (2024a) as a Web agent explores vision-language models, ranking models, and a bounding box annotation tool to improve Web elements grounding from lengthy and noisy HTML codes. Consequently, using tools in a multi-agent system enhances the invariance and independence from irrelevant contexts, ensuring that their operations are streamlined and focused solely on necessary information.
4.4 Towards Orderability of Preferences & Invariance & Independence from Irrelevant Context through Neuro-Symbolic Reasoning
Alignment with Dual-Process Theories
Neural-symbolic reasoning is another promising approach to achieving consistent ordering of preferences and invariance by combining the strengths of languages and symbolic logic in a multi-agent system. Nye et al. (2021) bridge neural-symbolic reasoning with the dual-process theories. It posits that LLMs are predominantly single systems operating in a mode akin to System 1. To address this, it introduces a multi-agent system that includes System 2-like agents to break down complex tasks into symbolic programs, filter candidate responses generated the previous LLM, and compare them with a minimal world model to check consistency. Such a slower interaction between LLMs and neuro-symbolic modules enhances the coherence and explainablity of the overall reasoning process.
Coherent Orderability of Preference
A multi-agent system incorporating symbolic modules can not only understand language queries but also solve them with a level of consistency, providing a faithful and transparent reasoning process based on well-defined rules that adhere to logical principles, which is unachievable by LLMs alone. Logic-LM Pan et al. (2023) combines problem formulating, symbolic reasoning, and result interpreting agents, where the symbolic reasoner empowers LLMs with deterministic symbolic solvers to perform inference, ensuring a correct answer is consistently chosen. Its multi-agent framework also encourages self-refinement that modifies logical formulation errors using error messages from the symbolic reasoner as the feedback. Besides, SymbolicToM Sclar et al. (2023b) and KRISP Marino et al. (2021) construct explicit symbolic graphs and answer questions by retrieving nodes in the graph. Binder Cheng et al. (2022), Parsel Zelikman et al. (2023), LEFT Hsu et al. (2024), and Fang et al. (2024) decompose tasks into planning, parsing, and execution, where the symbolic reasoning agents can help maintain a coherent order of preferences among symbolic options in the system outputs. By skipping the symbolic module, Parsel Zelikman et al. (2023) observes a substantial performance drop of $19.5\%$. LEFT Hsu et al. (2024) also outperforms end-to-end baselines without symbolic programs by $3.85\%$ on average across multiple experiments. In more explicit scenarios, logical modules can directly compare the order of options represented as variables—such as “left" or “right" in relational logic Hsu et al. (2024)—rather than relying on a single LLM to generate responses indeterministically within the natural language space.
Abstraction that Boils Down to Logical Essence
Neuro-Symbolic modules typically expect standardized input formats Zelikman et al. (2023); Pan et al. (2023); Sclar et al. (2023b); Hsu et al. (2024); Fang et al. (2024); Yang et al. (2024); Subramanian et al. (2024), analogous to making API calls with external tools. This layer of abstraction enhances the independence from irrelevant contexts and maintains the invariance of LLMs when handling natural language queries. The only relevant factor is the parsed inputs into the predetermined neuro-symbolic programs. For instance, Ada Wong et al. (2023) introduces symbolic operators to abstract actions, ensuring that lower-level planning models are not compromised by irrelevant information in the queries and observations. Without the symbolic action library, a single LLM would frequently fail at grounding objects or obeying environmental conditions, resulting in a significant accuracy gap of approximately $59.0$-$89.0\%$.
4.5 Towards Orderability of Preferences & Invariance through Reflection, Debate, and Memory
Alignment with Dual-Process Theories
Due to the probabilistic outputs of LLMs, which resemble the rapid, non-iterative nature of human System 1 cognition, ensuring preference orderability and invariance is challenging. In contrast, algorithms that enable self-reflection and multi-agent systems that promote debate and consensus can help align outputs more closely with the deliberate and logical decision-making typical of System 2 processes, thus enhancing rational reasoning in agents.
Memory is one of the most fundamental cognitive processes that lead to reasoning, creativity, learning, and even self-consciousness in humans Solso and Kagan (1979); Craik and Lockhart (1972); Leydesdorff and Hodgkin (2017); Johnson-Laird (1983); Laird (2019); Sun (2001). In other words, even an artificial general intelligence (AGI) that lacks memory and forgets every past conversation after each query would find it hard to develop a coherent and rational decision-making.
Deliberation that Slows Down the “Thinking" Process
A narrow definition of the agent memory includes historical information within the same conversation Zhang et al. (2024b), like the multi-round self-reflection prompting strategies that encourage agents to critically evaluate their previous responses Shinn et al. (2024); Madaan et al. (2024); Wang et al. (2022b); Zhong et al. (2024); Lu et al. (2023). Historical conservations, usually from different or opposite perspectives that corrects initial errors, help produce a more error-prone and consistent final decision. Yao et al. (2024, 2023b); Besta et al. (2024) even support backtracking to further slow down the “thinking" process.
A broader definition of agent memory expands to historical information across multi-tasks and multi-agents Zhang et al. (2024b). Corex Sun et al. (2023a) finds that orchestrating multiple agents to work together yields better complex reasoning results, exceeding strong single-agent baselines Wang et al. (2022b) by an average of $1.1$-$10.6\%$. Retroformer Yao et al. (2023a) equips the single-agent Reflexion Shinn et al. (2024) algorithm with an additional LLM to generate verbal reinforcement cues and assist its self-improvement, enhancing accuracy by $1.0$-$20.9\%$. ChatEval Chan et al. (2023) introduces a multi-agent debate framework to mimic human annotators collaborating in robust answer evaluations. Its multi-agent approach achieves greater alignment with human preferences compared to single-agent evaluations, enhancing accuracy by $6.2\%$ for GPT-3.5 and $2.5\%$ for GPT-4, and an increase of $16.3\%$ and $10.0\%$ in average Spearman and Kendall-Tau correlations Zhong et al. (2022) with human judgements in GPT-4. MetaAgents Li et al. (2023j) effectively coordinates agents within task-oriented social contexts to achieve consistent behavior patterns, and the implementation of agent reflection in this system leads to a $21.0\%$ improvement in success rates.
LM vs LM Cohen et al. (2023), FORD Xiong et al. (2023), Multi-Agent Debate Liang et al. (2023); Du et al. (2023), DyLAN Liu et al. (2023d), and Khan et al. (2024) highlight the profound impact of multi-agent collaboration through cross-examination and debates. These studies demonstrate substantial improvements in performance when multiple agents are orchestrated to work in collaboration. Specifically, LM vs LM Cohen et al. (2023) illustrates how its multi-agent framework improves F1 scores by an average of $15.7\%$ compared to the single-agent baseline Yoshikawa and Okazaki (2023). FORD Xiong et al. (2023) reports an accuracy increase up to $4.9\%$ compared to a single LLM. Liang et al. (2023) indicates significant improvements in accuracy — $17.0\%$ for translation tasks and $16.0\%$ for reasoning tasks — by employing a multi-agent strategy, effectively bridging the performance gap between GPT-3.5 and GPT-4 by harnessing multi-agents. Du et al. (2023) finds that multi-agent debates not only enhance reasoning performance by $8.0$-$14.8\%$, but more importantly, increase factual accuracy by $7.2$-$15.9\%$. DyLAN Liu et al. (2023d) observes $3.5$-$4.1\%$ in accuracy improvements over single-agent execution. Multi-agent debating in Khan et al. (2024) also leads to more truthful answers, boosting single-agent baselines by $28.0\%$. Multi-Agent Collaboration Talebirad and Nadiri (2023), ChatDev Qian et al. (2023), AgentCF Zhang et al. (2023), AutoGen Wu et al. (2023), Social Learning Mohtashami et al. (2023), S3 Gao et al. (2023a), Ke et al. (2024), and Chern et al. (2024) continue to push the frontier of a multi-agent system’s applications beyond daily conversation to a versatile set of real-world task completions.
LLMs are also sensitive to prompt perturbations due to token bias and noises Sclar et al. (2023a). One of the most worrying examples are adversarial attacks Gehman et al. (2020); Ganguli et al. (2022); Du et al. (2022); Wei et al. (2024); Perez et al. (2022); Zou et al. (2023) through malicious prompt engineering, also named the Red Team Task. Chern et al. (2024) introduce a multi-agent debating approach with agents having harmless, neutral, or harmful agent intentions. It finds that multi-round multi-agent debate surpasses the self-reflection of a single agent, thus improving the invariance.
These collaborative approaches, in summary, allow each agent in a system to compare and rank its preference on choices from its own or from other agents through critical judgments. It helps enable the system to discern and output the most dominant decision as a consensus, thereby improving the orderability of preference. At the same time, through such a slow and critical thinking process, errors in initial responses or input prompts are more likely to be detected and corrected. Accumulated experience from past error planning contributes to a self-evolving process within the multi-agent system Zhang et al. (2024b), resulting in a final response or a consensus that is less sensitive to specific wording or token bias, moving the response towards better consistency and invariance.
5 Evaluating Rationality of Agents
The amount of studies for testing rationality in multi-modal and multi-agent systems remains scant, despite the growing interest in the field. While there are numerous reasoning benchmarks available, including commonsense reasoning (Talmor et al., 2019), logical reasoning Liu et al. (2021, 2023a), multi-hop reasoning (Yang et al., 2018), mathematical reasoning (Hendrycks et al., 2021), structured data reasoning Chen et al. (2020), conceptual reasoning Zhou et al. (2024), and general capabilities through multi-agent evaluations Rasheed et al. (2024); Ma et al. (2024); Wang et al. (2024b); Abdelnabi et al. (2023).
However, they do not directly measure rationality. Many of these benchmarks fail to prove whether reasoning is actually used in solving the tasks, leaving no guarantee that these tasks will be solved consistently when generalized to other representations or domains. Issues such as data contamination Magar and Schwartz (2022); Dong et al. (2024); Sainz et al. (2023); Jacovi et al. (2023) further compound the problem, as some benchmarks may inadvertently include the training data of these LLMs, leading to inflated performance scores. Hence, even though solid reasoning will imply rationality, existing approaches fall short in making the logic click. In this section, we point to several existing ingredients that can constitute the bread-and-butter of future generations of evaluation approaches for rationality.
Adapting Cognitive Psychology Experiments
Recent works propose adapting vignette-based experiments borrowed from cognitive psychology to test whether LLMs are susceptible to cognitive biases and fallacies. For instance, Binz and Schulz (2023) tested GPT-3 on the conjunction fallacy, finding that they exhibit human-like biases. However, many of these approaches are informal and subjective, failing to scale in a way that allows for drawing statistically significant conclusions. Moreover, LLMs may be subject to cognitive biases not existent in humans, such as the hypothetical "algorithmic bias" proposed by Bender et al. (2021), which could lead to unintended consequences in decision-making tasks. Further research is needed to uncover and characterize these potential biases.
Testing Grounding against Hallucination
Information grounding is usually evaluated by the level of hallucination Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023), which can impact the rationality of agent systems. Multiple evaluation benchmarks targeting language-only dialogue have been proposed, such as BEGIN Dziri et al. (2022b), HaluEval Li et al. (2023f), DialFact Gupta et al. (2021), FaithDial Dziri et al. (2022a), AIS Rashkin et al. (2023), and others Zheng et al. (2023b); Das et al. (2023); Cao et al. (2021). In contrast, benchmarks on multi-agent frameworks beyond language dialogue or those involving multi-modalities are very limited. Liu et al. (2024a) moves beyond conversation to code generation; EureQA Li et al. (2023a) focuses on reasoning chains; and TofuEval Tang et al. (2024) evaluates hallucination in multi-domain summarization. Object hallucination Rohrbach et al. (2018); Biten et al. (2022), POPE Li et al. (2023h), and LLaVA-RLHF Sun et al. (2023b) are the few examples evaluating multi-modal hallucination. The community needs more hallucination benchmarks to quantitatively evaluate the extent to which multi-modal and multi-agents reduce hallucinations in comparison with baselines.
Testing the Orderability of Preference
There are almost no benchmarks for evaluating whether LLMs or agents have a consistent preference in the selection of available options. The Multiple Choice Problem (MCP) serves as a common testing ground. Zheng et al. (2023a) shows that LLMs are susceptible to changes in the positioning of options. Since the underlying logic remains the same, it also makes LLMs fail to pass the property of invariance. There are tons of MCP benchmarks PaperswithcodeMCQA , but solely on the accuracy of selections, overlooking the consistency of preference. However, Robinson et al. (2023) highlights that the Proportion of Plurality Agreement (PPA) offers a measure of order invariance that does not depend on the model’s ability to perform a task, suggesting a promising direction. We also need evaluations into vision and other modalities.
Testing the Principle of Invariance
The realization of language can be independent of its meaning Frege (1892); Wittgenstein (1953); Heineman (2023). Recent studies within the realm of data contamination investigate whether LLMs can generate consistent responses across different, yet inherently equivalent, framings of the same task. These studies introduce perturbations to the original task descriptions to assess whether LLMs’ responses will change significantly. Perturbation techniques include modifying instruction templates Weber et al. (2023), paraphrasing task descriptions (Yang et al., 2023; Ohmer et al., 2024), or altering the order of in-context learning exemplars Lu et al. (2021); Pecher et al. (2024). Specifically, approaches involve some versions of paraphrasing or permutation, such as changing the instruction templates Weber et al. (2023), rewording task descriptions (Yang et al., 2023; Ohmer et al., 2024; Wang et al., 2024b), translating the prompts into a different language (Ohmer et al., 2023, 2024; Xu et al., 2024a) and then back to the original language Yang et al. (2023), and making subtle changes to entities in task descriptions without affecting the logical structure, like altering names of the characters, numerical values in math problems, or locations of the events Wang et al. (2024b). Permutation also includes reordering in-context learning examples Lu et al. (2021); Pecher et al. (2024) and, in the case of multiple-choice questions, rearranging the options (Zong et al., 2023; Zheng et al., 2023a).
It is crucial to recognize that these perturbations are superficial; the altered task descriptions remain syntactically and semantically equivalent to their originals, although linguistic expressions or narratives may vary substantially. This observation highlights the need for developing methods that go beyond surface-level perturbations to effectively evaluate the robustness and invaraince of LLMs across diverse problem framings and modalities.
Testing Independence from Irrelevant Context
We need more expansive perturbations to evaluate the independence from irrelevant context. Studies such as Shi et al. (2023), Wu et al. (2024), Liu et al. (2024d), and Yoran et al. (2023) have explored the phenomenon of “lost-in-context" by introducing random or misleading sentences into original problem statements. While earlier benchmarks like those from Weston et al. (2015), Sinha et al. (2019), Clark et al. (2020), and Webson and Pavlick (2021) have included irrelevant content, they have been predominantly limited to language modalities and single-agent systems. Recent benchmarks such as MileBench Song et al. (2024), Mementos Wang et al. (2024c), Seed-bench-2 Li et al. (2023b), and DEMON Li et al. (2023d) begin to evaluate multi-modal agents in long context or image sequences, where accurately responding to a specific question requires isolating only the relevant information from the long context window.
6 Open Problems and Future Directions
Inherent Rationality
It is important to understand that integrating most of these agents or modules with LLMs still does not inherently make LLMs more rational. Current methods are neither sufficient nor necessary, but they serves as instrumental tools that bridges the gap between an LLM’s response and rationality. These approaches enables multi-agent systems, which are black boxes from the user’s perspective, to more closely mimic rational thinking in their output responses. However, despite these more rational responses elicited from multi-modal and multi-agent systems, the challenge of how to effectively close the loop and bake these enhanced outputs back into the LLMs Zhao et al. (2024), beyond mere fine-tuning, remains an open topic. In other words, can we leverage these more rational outputs to inherently enhance a single foundation model’s rationality in its initial responses in future applications?
More Comprehensive Evaluation on Rationality
Section 4 thoroughly compares multi-modal and multi-agent systems over their LLM-based single-agent baselines. However, the choices of evaluation metrics are important Schaeffer et al. (2024); these examples predominantly focus on the accuracy of their final performance, ignoring the most interesting intermediate reasoning steps and the concept of rationality. Section 5 furthermore acknowledges that while there have been some efforts to assess the rationality of agent systems, the field still lacks comprehensive and rigorous evaluation metrics. Moreover, most existing benchmarks on rationality provide limited comparisons between multi-agent frameworks and single-agent baselines, thus failing to fully elucidate the advantages multi-agent frameworks can offer.
Future research should prioritize the development of more robust and scalable methods for evaluating rationality, taking into account unique challenges and biases posed by agents. A promising direction is to create benchmarks specifically tailored to assess rationality, going beyond existing ones on accuracy. These new benchmarks should avoid data contamination and emphasize tasks that demand consistent reasoning across diverse representations and domains. There is a need for more rigorous and large-scale studies on the principles of invariance and orderability of preference, together with their applications to testing rationality in agent systems. This would involve developing more sophisticated perturbation methods that probe the consistency of reasoning at a deeper level, as well as designing experiments that yield statistically significant results.
Encouraging More Multi-Modal Agents in Multi-Agent Systems
Research into the integration of multi-modality within multi-agent systems would be promising. Fields such as multi-agent debate, collaboration, and neuro-symbolic reasoning, as shown in Figure 1, currently under-utilize the potential of multi-modal sensory inputs. We believe that expanding the role of multi-modalities, including but not limited to vision, sounds, and structured data could significantly enhance the capabilities and rationality of multi-agent systems.
7 Conclusions
This survey builds connections between multi-modal and multi-agent systems with rationality, guided by dual-process theories and the four axioms we expect a rational agent or agent systems should satisfy: information grounding, orderability of preference, independence from irrelevant context, and invariance across logically equivalent representations. Our findings suggest that the grounding can usually be enhanced by multi-modalities, world models, knowledge retrieval, and tool utilization. The remaining three axioms are typically intertwined, and we sometimes describe their collective characteristics informally using terms such as coherence, consistency, and trustworthiness. These axioms are simultaneously improved by achievements in multi-modalities, tool utilization, neuro-symbolic reasoning, self-reflection, and multi-agent collaborations. These fields of research, by either slowing down the “thinking" process or boiling down real-world problems to logical essence, mimic the “System 2" thinking in human cognition, thereby enhancing the rationality of multi-agent systems in decision-making scenarios, compared to single-agent language-only baselines that resemble the “System 1" process.
Collaboration between the AI research community and cognitive psychologists could be particularly fruitful. We need better evaluation benchmarks on the rationality of agents, more exploration to mitigate cognitive biases in multi-modal and multi-agent systems, and deeper understanding of how these biases arise and how they can be mitigated, ultimately enhancing rationality in decision-making processes.
8 Limitations
The fields of multi-modal and multi-agent systems are rapidly evolving. Despite our best efforts, it is inherently impossible to encompass all related works within the scope of this survey. Our discussion also possesses limited mention of the reasoning capabilities, theory of mind in machine psychology, and cognitive architectures, all of which lie beyond the scope of this survey but are crucial for a deeper understanding of LLMs and agent systems. Furthermore, the concept of rationality in human cognitive science may encompass more principles and axioms than those defined in our survey.
Appendix A Orderability of Preferences.
Comparability
When faced with any two alternatives A and B, the agent should have at least a weak preference, i.e., $A\succeq B$ or $B\succeq A$. This means that the agent can compare any pair of alternatives and determine which one is preferred or if they are equally preferred.
Transitivity
If the agent prefers A to B and B to C, then the agent must prefer A to C. This ensures that the agent’s preferences are consistent and logical across multiple comparisons.
Closure
If A and B are in the alternative set S, then any probabilistic combination of A and B (denoted as ApB) should also be in S. This principle ensures that the set of alternatives is closed under probability mixtures.
Distribution of probabilities across alternatives
If A and B are in S, then the probability mixture of (ApB) and B, denoted as [(ApB)qB], should be indifferent to the probability mixture of A and B, denoted as (ApqB). This principle ensures consistency in the agent’s preferences when dealing with probability mixtures of alternatives.
Solvability
When faced with three alternatives A, B, and C, with the preference order $A\succeq B\succeq C$, there should be some probabilistic way of combining A and C such that the agent is indifferent between choosing B or this combination. In other words, the agent should be able to find a solution to the decision problem by making trade-offs between alternatives.
One consequence of the orderability is the concept of dominance: If alternative A is better than alternative B in terms of one attribute and at least as good in terms of all other attributes, the dominant option A should be chosen. An example of a fallacy that violates dominance is the sunk cost fallacy, where an agent continues to invest in a suboptimal alternative due to past investments, despite the availability of better options based on future outcomes.
Appendix B The Orderability of Preferences Matters for LLM-based Evaluations
This section talks about LLM-based evaluation rather than evaluating the rationality of LLMs discussed in Section 5. Recent research underscores a critical need for more rational LLM-based evaluation methods, particularly for assessing open-ended language responses. CoBBLEr Koo et al. (2023) provides a cognitive bias benchmark for evaluating LLMs as evaluators, revealing a preference for their own outputs over those from other LLMs. Stureborg et al. (2024) argues that LLMs are biased evaluators towards more familiar tokens and previous predictions, and exhibit strong self-inconsistency in the score distribution. Luo et al. (2023); Shen et al. (2023); Gao et al. (2023c); Wang et al. (2023b); Chen et al. (2023); Chiang and Lee (2023); Zheng et al. (2024b); Fu et al. (2023); Liu et al. (2023c) also point out the problem with a single LLM as the evaluator, with concerns over factual and rating inconsistencies, a high dependency on prompt design, a low correlation with human evaluations, and struggles with the comparison. As a result, having a coherent orderability of preferences aligned with human preference becomes increasingly important.
Multi-agent systems might be a possible remedy. By involving multiple evaluative agents from diverse perspectives, it becomes possible to achieve a more balanced and consistent orderability of preferences. For instance, ChatEval Chan et al. (2023) posits that a multi-agent debate evaluation usually offers judgments that are better aligned with human annotators compared to single-agent ones. Bai et al. (2024) also finds decentralized methods yield fairer evaluation results.